# Rethinking Softmax: Self-Attention with Polynomial Activations

### v1

@@ 5 Experiments
In this section, we validate the theory from section 4 on a variety of transformer tasks. We perform the empirical validation on two primary activations from section 4, namely a cubic polynomial activation $\phi(x)=x^{3}$ and a linear polynomial $\phi(x)=x$. The goal will be to show that by suitably scaling these activations using the theory in section 4, we can achieve competitive performance when compared to softmax. For the rest of this section we will simply denote these activations by $x^{3}$ and $x$.
Figure 1: Training ViT-Tiny with the activation $\phi(x)=x^{3}$ with different sequence lengths and different scales. As the sequence length gets smaller, the log scale needed to obtain good accuracy decreases validating the theory from section 4.2.
@@ 5.1 Image classification
5.1.1 ViT-Tiny on Tiny-Imagenet:
In this section we test the theory from section 4 on the ViT-Tiny architecture (Steiner et al., 2021) trained from scratch on the Tiny-Imagenet dataset (Le & Yang, 2015).
Our first experiment was to test how the Top-$1\%$ accuracy changes for a ViT-Tiny trained on Tiny-Imagenet as we change the sequence length of the input and the scale predicted in corollaries 4.3 and 4.5 when using the activation $x^{3}$. According to the theory developed in section 4, the Frobenius norm scales according to $\mathcal{O}(\sqrt{N})$ when we scale $X^{3}$ by $\frac{1}{\sqrt{N}}$. Thus as the sequence length decreases we should see the amount of scaling in a Log scale decrease.
Figure 1 shows the results of this experiment. We considered four different input sequence lengths of sizes $256$, $64$, $16$ and $8$. We ran several ViT-Tiny architectures with a variety of scalings of the form $\mathcal{O}(\frac{1}{\sqrt{N}})$ where $N$ ranged below to above the sequence length. As can be seen from figure 1 as the sequence length got smaller the amount of scaling, shown in Log scale on the x-axis, needed for good accuracy got smaller verifying the theory in section 4.2.
The second experiment compared activations $x^{3}$ and $x$, along with scaled versions $\frac{1}{16}x^{3}$ and $\frac{1}{16}x$, against softmax using the Tiny-ViT architecture on Tiny-Imagenet. With a sequence length of $256$ ($\sqrt{256}=16$), we decided to take $\frac{1}{16}$ as the scale of the polynomial activations. The experiment used a patch size of 4, 3 attention heads, and 12 layers as described in Steiner et al. (2021). Results in table 1 show $\frac{1}{8}x^{3}$ outperforming softmax, while the unscaled version performed poorly. Similarly, $\frac{1}{16}x$ performed competitively with a significant drop in performance without scaling.
Figure 2 displays the Frobenius norm of the self-attention matrix during training for five activations in layers 2 and 12 of ViT-Tiny, averaged across all heads. Norms for $x^{3}$ and $x$ are higher than softmax, but scaling by $\frac{1}{16}$ reduces them to more stable levels, improving training stability. Similarly, figure 3 shows the Jacobian’s Frobenius norm, where scaling also brings the norms closer to softmax, ensuring more stable gradients. Further plots for other layers are in appendix A.2.2.
softmax $\frac{x^{3}}{16}$ $x^{3}$ $\frac{x}{16}$ $x$
Top-1% accuracy 50.26 50.5 45.3 47.9 31.78
Table 1: Comparison of Top-1% accuracy on Tiny-Imagenet between softmax and polynomial activations. The cubic activation outperforms softmax when the right scale of $\frac{1}{8}$ is applied. Similarly, the linear activation is competitive only with an optimal scale.
Figure 2: Frobenius norm of the self-attention matrix with five different activations in layer 2 and 12 of the ViT-Tiny architecture during training.
Figure 3: Frobenius norm of the Jacobian of the self-attention matrix with five different activations in layer 2 and 12 of the ViT-Tiny architecture during training.
5.1.2 Larger vision transformers on ImageNet-1k
For this experiment we carried out an image classification task using a variety of different vision transformers from the literature on the ImageNet-1k dataset. We found that the scales $\frac{1}{14}$ worked best for both $x^{3}$ and $x$.
We train all models on the ImageNet-1k dataset from scratch and report Top-1 accuracy on the validation set. We use PyTorch Paszke et al. (2019) and Timm Wightman (2019) libraries to train our models with similar setups to He et al. (2022) Liu et al. (2021). We examined our approach along with the following three transformer architectures to show its generalization:
- ViT: Dosovitskiy et al. (2020) is the pioneering work that interprets an image as a sequence of patches and processes it by a standard Transformer encoder as used in NLP. This simple, yet scalable, strategy works surprisingly well when coupled with pre-training on large datasets. We use ViT-Small which has the following settings: patch size = 16, embedding dimensions = 384, number of heads = 6, and layers = 12. Also, We use ViT-Base which has the following settings: patch size = 16, embedding dimensions = 768, number of heads = 12, and layers = 12.
- DeiT: Touvron et al. (2021) is a well-known transformer based on ViT. It is very similar to ViT except that it converged faster due to better training recipe since we do not use the distillation token imposed in DeiT. We use DeiT-Small which has the following settings: patch size = 16, embedding dimensions = 384, number of heads = 6, and layers = 12. Also, We use Deit-Base which has the following settings: patch size = 16, embedding dimensions = 768, number of heads = 12, and layers = 12.
- Swin Transformer: Liu et al. (2021) produces a hierarchical feature representation and proposes the shifted window based self-attention which is shown to be effective and efficient on vision problems. We use Swin-Small with 96 channels and Swin-Base with 128 channels in the hidden layers of the first stage. The window size is set to $M$ = 7 by default, the query dimension of each head is $d$ = 32, and layer numbers are ${2,2,18,2}$ for all experiments.
- XciT: Xiong et al. (2021) is a vision transformer architecture consisting of two different components compared to the standard ViT. Firstly, it has Local Patch Interaction in each block, which includes one depth-wise 3×3 convolution followed by Batch Normalization, GELU, and another depth-wise 3×3 convolution. Secondly, it uses Cross-Covariance attention, where the attention map is derived from the cross-covariance matrix computed over the key and query projections of the token features. We use XCiT-S12 with a patch size of 16 and XCiT-M24 with a patch size of 24.
The results are shown in table 2. The activation $\frac{1}{14}x^{3}$ performed best on the ViT’s and the Swin transformers while softmax performed best on the DeiT architectures. Further ablations with different scales and activations can be fond in appendix A.2.1.
Activation Models
ViT-Base ViT-Small DeiT-Base DeiT-Small Swin-Base Swin-Small XciT-Medium XciT-Small
softmax 79.6 80.2 78.9 79.6 83.0 83.3 81.1 81.2
$\frac{x^{3}}{14}$ 79.6 80.5 77.4 78.3 83.2 83.4 81.2 82.1
$x^{3}$ 77.8 78.6 76.1 76.3 79.7 79.9 78.1 78.3
$\frac{x}{14}$ 76.9 77.8 79.6 79.8 79.4 79.5 79.2 79.3
$x$ 73.2 73.9 77.7 77.9 77.8 77.9 76.4 76.6
@@ 5.1.2 Larger vision transformers on ImageNet-1k
Table 2: Comparsions of pre-training models with different activation functions on ImageNet-1k datasets. We report the classification top-1 accuracy (%).
Figure 4 plots the Frobenius norm of the self-attention matrix with the five different activations during training in layers 2 and 12, averaged over all heads within each layer, of the ViT-Small architecture. By scaling the activations $x^{3}$ and $x$ by $\frac{1}{14}$ we were able to control the scale of the Frobenius norm of the self-attention matrix and obtain scales comparable to softmax’s scale. Simiarly, figure 5 plots the Frobenius norm of the Jacobian of the self-attention matrix during training for layers 2 and 12, averaged over all heads. By scaling the activations $x^{3}$ and $x$ by $\frac{1}{14}$ we were able to control the scale of the Jacobian norm and obtain scales comparable to softmax’s scale during training. Plots for other layers of the architecture during training are given in appendix A.2.2.
Figure 4: Frobenius norm of the self-attention matrix with five different activations in layer 2 and 12 of the ViT-Tiny architecture during training.
Figure 5: Frobenius norm of the Jacobian of the self-attention matrix with five different activations in layer 2 and 12 of the ViT-Tiny architecture during training.
5.1.3 Visualizing self-attention with ViT-Base
We plotted heat maps of self-attention matrices using the $\frac{1}{14}x^{3}$ and softmax activations across two layers and heads after convergence, averaging over a fixed training batch of size 128. Figure 6 shows layer 2, head 8, highlighting the differences in attention patterns, with $\frac{1}{14}x^{3}$ containing both positive and negative values. Similarly, Figure 7 for layer 12, head 6 shows distinct patterns for each activation. Overall, the ViT-Base architecture with $\frac{1}{14}x^{3}$ exhibits notably different self-attention patterns compared to softmax.
We visualized how the self-attention matrix targets different regions of an image. Using an image from the ImageNet-1k validation set, we extracted the class token, reshaped it into a $(14,14)$ grid representing 196 patches, and then mapped it back to the original image size with nearest neighbor interpolation. Figure 8 shows the input image, while figure 9 illustrates the self-attention matrices for different activations in layer 12, head 6 of the ViT-Base architecture after convergence, highlighting their distinct focus areas.
@@ 5.1.3 Visualizing self-attention with ViT-Base
Figure 6: Heat maps of the self-attention matrix in layer 2, head 8, of a ViT base architecture, comparing $\frac{1}{14}x^{3}$ (left) and softmax (right) activations after training. The stark difference in self-attention patterns between the two activations is evident, showing distinct distributions across input tokens.
@@
Figure 7: Heat maps of the self-attention matrix in layer 12, head 6, of a ViT base architecture, comparing $\frac{1}{14}x^{3}$ (left) and softmax (right) activations after training. The contrast in self-attention patterns between the two activations is clearly visible.
Figure 8: Fish image from validation set of ImageNet-1k broken up into $196=14\times 14$ patches.
Figure 9: Comparing how $\frac{1}{14}x^{3}$ and softmax self-attention matrices focus on different parts of the fish image from figure 8 in layer 12, head 6, after the model has converged.
@@ 5.2 Object Detection and Instance Segmentation
In this section, in order to examine the transfer learning ability of our models, we demonstrate our approach to object detection and segmentation tasks by fine-tuning our ImageNet-pretrained XCiT model on them. Our experiments are conducted on COCO 2017Lin et al. (2014), which has 118K training images and 5K validation images with 80 categories. We integrate the XCiT architecture as the backbone in the Mask R-CNN (He et al., 2017) detector with a Feature Pyramid Network (FPN). Due to XCiT’s inherently columnar design, we adapt it for FPN compatibility by extracting features from various layers for XCiT-S12. These features have a consistent stride of either 16. The feature resolutions are then adjusted to strides of [4, 8, 16, 32] This downsampling is accomplished through max pooling, while upsampling is achieved using a single transposed convolution layer. The model is trained for 36 epochs using the AdamW optimizer with learning rate of $10^{-4}$, 0.05 weight decay and 16 batch size. In table 3,we condunct experiments on XCiT-S12 models using 16×16 patches with the activations $\frac{1}{14}x^{3}$, $\frac{1}{14}x$ and softmax. We found we couldn’t train with the activations $x^{3}$ and $x$ on this task well so only report the others.
Activation $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
@@
$\frac{x^{3}}{14}$ 44.8 66.3 49 40.1 63.1 42.9
$\frac{x}{14}$ 44.8 66.2 49 40.1 63.1 42.8
Table 3: COCO object detection and instance segmentation performance on the mini-val set. All backbones are pretrained on ImageNet-1k, and use Mask R-CNN model. $AP^{b}$: Average Precision for bounding box predictions, $AP^{b}_{50/75}$: Average Precision at an Intersection over Union (IoU) threshold of 0.50/0.75 for bounding box predictions, $AP^{m}$: Average Precision for mask predictions, $AP^{m}_{50/75}$: Average Precision at an Intersection over Union (IoU) threshold of 0.50/0.75 for mask predictions
5.3 Natural language processing(NLP)
To assess the effectiveness of our approach on NLP tasks, we trained models on five benchmarks from the Long Range Arena (LRA) suite Tay et al. (2020): ListOps, Text Classification, Retrieval, Image Classification, and Pathfinder. We evaluated the activations $\frac{x^{3}}{14}$ and $\frac{x}{14}$ against softmax, finding that $x^{3}$ and $x$ did not train effectively on their own, so only results for these scaled activations and softmax are presented. Our implementation followed the guidelines from Xiong et al. (2021). The results are summarized in table 4.
Activation ListOps Text Retrieval Image Pathfinder
@@ 5.3 Natural language processing(NLP)
$\frac{x^{3}}{14}$ 37.5 63.3 80.9 37.2 68.7
$\frac{x}{14}$ 34.3 62.9 81.5 39.0 69.1
Table 4: Comparsions of transformer models with different activation functions on NLP tasks. We report the accuracy (%) on LRA benchmarks.
@@ 6 Limitations
While our work introduces novel activations that challenge the conventional softmax approach, there are some limitations to address. Our theoretical framework is primarily designed for dot-product self-attention and may not immediately extend to other attention mechanisms, although our empirical results showed competitive performance against softmax across different architectures. Additionally, we observed that while our activations performed well on vision tasks, their performance was less consistent on NLP tasks, suggesting that a more refined theoretical approach may be needed for these applications.
@@ 7 Conclusion
This work challenges the traditional view that transformer activations for attention must produce sparse probability distributions. We introduced a theoretical framework analyzing the Frobenius norm of the self-attention matrix, which suggests key scaling properties for activations in attention mechanisms. We proved that specific polynomial activations, which behave very differently from softmax, satisfy these properties. Through extensive experiments across vision and NLP tasks, we demonstrated that these alternative activations not only compete with but sometimes outperform softmax, offering a fresh perspective on attention mechanisms in transformers.
@@ A.1 Theoretical analysis
$\mathbf{softmax}\bigg{(}\begin{bmatrix}a_{11}&\cdots&a_{1n}\\ \vdots&\vdots&\vdots\\ a_{n1}&\cdots&a_{nn}\end{bmatrix}\bigg{)}=\begin{bmatrix}\frac{e^{a_{11}}}{\sum_{j=1}^{n}e^{a_{1j}}}&\cdots&\frac{e^{a_{1n}}}{\sum_{j=1}^{n}e^{a_{1j}}}\\ \vdots&\vdots&\vdots\\ \frac{e^{a_{n1}}}{\sum_{j=1}^{n}e^{a_{nj}}}&\cdots&\frac{e^{a_{nn}}}{\sum_{j=1}^{n}e^{a_{nj}}}.\end{bmatrix}$ (A.1)
@@
$||\mathbf{softmax}(A)||_{F}^{2}$ $=\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{1j}}}\bigg{)}^{2}(e^{2a_{11}}+\cdots e^{2a_{11}})+\cdots+\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{Nj}}}\bigg{)}^{2}(e^{2a_{N1}}+\cdots e^{2a_{NN}})$ (A.2)
$\leq\bigg{[}\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{1j}}}\bigg{)}(e^{a_{11}}+\cdots e^{a_{11}})\bigg{]}^{2}+\cdots+\bigg{[}\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{Nj}}}\bigg{)}(e^{a_{N1}}+\cdots e^{a_{NN}})\bigg{]}^{2}$ (A.3)
$=1+\cdots+1$ (A.4)
$=N$ (A.5)
@@
We move on to prove the second inequality in the statement of theorem 4.1. For this, let us write each entry of the matrix on the right of equation A.1 as follows:
$F_{kl}=\frac{e^{a_{kl}}}{\sum_{j=1}^{N}e^{a_{kj}}}.$ (A.6)
@@
$\frac{\partial}{\partial x_{ij}}F_{ij}$ $=F_{ij}-F_{ij}^{2}$ (A.7)
$\frac{\partial}{\partial x_{ik}}F_{ij}$ $=-F_{ij}F_{ik}\text{ for any }k\neq j$ (A.8)
$\frac{\partial}{\partial x_{kl}}F_{ij}$ $=0\text{ for any }k\neq i\text{ and }l\neq j.$ (A.9)
@@
$\nabla\mathbf{softmax}(A)=\begin{bmatrix}\nabla F_{11}\\ \vdots\\ \nabla F_{1N}\\ \nabla F_{21}\\ \vdots\\ \vdots\\ \nabla F_{NN}\end{bmatrix}$ (A.10)
@@
$\nabla F_{ij}=\begin{bmatrix}-F_{ij}F_{i1}^{2}&-F_{ij}F_{i2}&\cdots&F_{ij}-F_{ij}^{2}&\cdots&F_{ij}F_{iN}.\end{bmatrix}$ (A.11)
@@
$||\nabla\mathbf{softmax}(A)||_{F}^{2}=||\nabla F_{11}||_{F}^{2}+\cdots+||\nabla F_{NN}||_{F}^{2}.$ (A.12)
@@
$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $=|F_{i1}-F_{i1}^{2}|^{2}+|F_{i1}F_{i2}|^{2}+\cdots+|F_{i1}F_{iN}|^{2}$ (A.13)
$\hskip 14.22636pt+|F_{i2}F_{i1}|^{2}+|F_{i2}-F_{i2}^{2}|^{2}+\cdots+|F_{i2}F_{iN}|^{2}$ (A.14)
@@
$\hskip 14.22636pt+|F_{iN}F_{i1}|^{2}+|F_{iN}F_{i2}|^{2}+\cdots+|F_{iN}-F_{iN}^{2}|^{2}$ (A.16)
$\leq(F_{i1})^{2}(|1-F_{i1}|+|F_{i2}|+\cdots+|F_{iN}|)^{2}$ (A.17)
$\hskip 14.22636pt(F_{i2})^{2}(|F_{i1}|+|1-F_{i2}|+\cdots+|F_{iN}|)^{2}$ (A.18)
@@
$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $\leq 4F_{i1}^{2}(\widehat{F_{i1}}+F_{i2}+\cdots+F_{iN})$ (A.21)
@@
$\hskip 14.22636pt+4F_{iN}^{2}(F_{i1}+F_{i2}+\cdots+\widehat{F_{iN}})$ (A.23)
$\leq 4(F_{i1}^{2}+\cdots F_{iN}^{2})$ (A.24)
$=4.$ (A.25)
@@
$||\nabla\mathbf{softmax}(A)||_{F}^{2}\leq 4N$ (A.26)
@@
$||\nabla\mathbf{softmax}(A)||_{F}\leq 2\sqrt{N}.$ (A.27)
@@
$\mathop{\mathbb{E}}(c_{11}^{2})$ $=\mathop{\mathbb{E}}((\sum_{i=1}^{D}a_{1i}b_{i1})^{2})$ (A.28)
@@
$\mathop{\mathbb{E}}(c_{11}^{2})$ $=\mathop{\mathbb{E}}((\sum_{i=1}^{D}a_{1i}b_{i1})^{2p})$ (A.29)

### v2

@@ 6 Experiments
In this section, we evaluate a simple polynomial activation as an alternative to the standard softmax across a range of transformer applications commonly studied in the literature. The goal is to empirically challenge the conventional softmax properties, positivity, row-normalization, and sparsity—and examine whether these conditions are truly necessary for good transformer performance.
Building on the theoretical foundations from Section 4, we focus on the cubic polynomial $x^{3}$ as a test case. Notably, $x^{3}$ introduces both positive and negative values, does not normalize rows to sum to 1, and generally produces dense attention matrices—violating all three traditional softmax conditions.
We consider two scaling strategies for $x^{3}$:
  • 1.
    Fixed scale: Following Section 4, we scale $x^{3}$ by the inverse square root of the sequence length (which is fixed throughout training), as theory suggests this maintains optimization stability.
  • 2.
    Dynamic scale: Recognizing that the assumption of i.i.d. normal-distributed $Q$, $K$, $V$, used in Section 4, holds primarily at initialization, we also explore a learnable scale. This dynamic scale is initialized as above but optimized during training.
Our experiments are not designed to achieve state-of-the-art results. Instead, they aim to question the softmax paradigm and show that alternative activations, even those violating softmax’s traditional properties, can still lead to effective transformer models. For each experiment we ran five trials and report the mean result.
@@ 6.1 Image Classification
We conducted an image classification task using various vision transformer architectures from the literature on the ImageNet-1k dataset. For this task, the standard sequence length employed by the vision transformers on the ImageNet-1k dataset is $196$.
We trained all models on the ImageNet-1k dataset from scratch and report Top-1 accuracy on the validation set. We examined our approach along with the following four transformer architectures to show its generalization, ViT-B [10], DeiT [25], Swin Transformer [17], XCiT [3]. Each transformer was trained following the approach in each paper. The results are shown in Table 2. As can be seen from the table the $x^{3}$ activation with a dynamic scale performed the best and the one with a fixed scale performed comparable to softmax. On the other hand $x^{3}$ (with no scale) underperformed on all ViTs showing how important the theory on scaling as developed in Section 4 is. For a discussion on other polynomials and linear attention see Section A.2.4.
@@
Table 2: Comparisons of pre-training models with different activation functions on ImageNet-1k. We report top-1 classification accuracy (%).
Models
ViT-B DeiT-B Swin-B XciT-M
softmax 80.3 81.5 83.5 82.7
$x^{3}$ + fixed scale 80.2 81.4 83.6 82.8
$x^{3}$ + dynamic scale 80.3 81.6 83.6 82.9
$x^{3}$ 78.1 78.5 79.9 79.5
Visualizing attention heads:
Softmax attention traditionally satisfies three key properties: positivity, row normalization, and sparsity. In contrast, the polynomial activation $x^{3}$ (with or without positive scaling) takes positive values for $x>0$ and negative values for $x<0$, thus violating these softmax constraints. To better understand how this affects attention patterns, we analyzed the self-attention matrices of ViT-B models trained with the $x^{3}$ + dynamic scale activation and compared them to those using softmax. We visualized heatmaps of the attention matrices after convergence, focusing on two representative layers and heads, and averaged over a fixed batch of 128 samples. Figure 4 shows results for layer 2, head 8, where the $x^{3}$ + dynamic scale activation produces attention scores with both positive and negative values, unlike softmax. Similarly, Figure 5 illustrates distinct patterns for layer 12, head 6, highlighting how the two activations differ in learned attention distributions.
Interpretability.
Even with these differences, the $x^{3}$ + dynamic scale activation still learns meaningful attention patterns. For instance, in Figure 4 (left), we observe 14 bands about the diagonal, indicating that the head has learned to attend to patches in the same image row as the query patch, sufficient for effective image classification. This suggests that the sign of attention values (positive or negative) is not inherently critical for attention allocation. Further, Figure 5 (left) reveals vertical lines, showing attention that depends solely on key positions, independent of the query position. Smaller key indices receive higher attention weights, focusing model capacity where it matters most for classification tasks. These findings demonstrate that, despite deviating from softmax properties, the $x^{3}$ + dynamic scale activation enables the model to discover effective attention patterns. We noticed similar attention patterns for the $x^{3}$ + fixed scale activation.
@@
Figure 4: Heat maps of the self-attention matrix in layer 2, head 8, of a ViT base architecture, comparing $x^{3}$ + dynamic (left) and softmax (right) activations after training. The stark difference in self-attention patterns between the two activations is evident, showing distinct distributions across input tokens.
@@
Figure 5: Heat maps of the self-attention matrix in layer 12, head 6, of a ViT base architecture, comparing $x^{3}$ + dynamic scale (left) and softmax (right) activations after training. The contrast in self-attention patterns between the two activations is clearly visible.
@@ 6.2 Object Detection and Instance Segmentation
In this section, in order to examine the transfer learning ability of our models, we demonstrate our approach on object detection and segmentation tasks by fine-tuning a ImageNet-pretrained XCiT model on them. Our experiments are conducted on COCO 2017 [16], which has 118K training images and 5K validation images with 80 categories. We integrate the XCiT architecture as the backbone in the Mask R-CNN [13] detector with a Feature Pyramid Network (FPN). Due to XCiT’s inherently columnar design, we adapt it for FPN compatibility by extracting features from various layers for XCiT-S12. These features have a consistent stride of 16. The feature resolutions are then adjusted to strides of [4, 8, 16, 32]. This downsampling is accomplished through max pooling, while upsampling is achieved using a single transposed convolution layer. We conducted experiments on XCiT-S12 models using 16×16 patches with the activations $x^{3}$ + fixed scale, $x^{3}$ + dynamic scale, with scale initialized to $\frac{1}{14}$, and softmax. We found when we trained with $x^{3}$ the model would not converge thereby suggesting that a scaling is extremely important as predicted by the theory in Section 4.2. The training hyperparameters are given in Section A.2.2. Table 3 shows the results of the experiment. From the table we see that the activation $x^{3}$ + dynamic scale obtains better performance to softmax with $x^{3}$ + fixed scale obtaining comparable performance.
Table 3: COCO object detection and instance segmentation performance on the mini-val set. All backbones are pretrained on ImageNet-1k, and use Mask R-CNN model. $AP^{b}$: Average Precision for bounding box predictions, $AP^{b}_{50/75}$: Average Precision at an IoU threshold of 0.50/0.75 for bounding box predictions, $AP^{m}$: Average Precision for mask predictions, $AP^{m}_{50/75}$: Average Precision at an IoU threshold of 0.50/0.75 for mask predictions.
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
@@
$x^{3}$ + fixed scale 44.8 66.3 49.1 40.2 63.1 43.0
$x^{3}$ + dynamic scale 45.1 66.5 49.4 40.4 63.2 43.1
6.3 Nyströmformer on LRA benchmark
Nyströmformer is a transformer designed to handle long-range dependencies more efficiently by approximating the self-attention mechanism with the Nyström method [30]. To assess the effectiveness of our approach, we trained models on five benchmarks from the Long Range Arena (LRA) suite [23]: ListOps, Text Classification, Retrieval, Image Classification, and Pathfinder.
On each dataset we compared a Nyströmformer with softmax, $x^{3}$ + dynamic scale, and $x^{3}$. For the $x^{3}$ + dynamic scale activation on each dataset we used a different scale as initialization based on the sequence length used in the original Nyströmformer model. For ListOps the original model used a sequence length of $512$ and so the scale we used as initialization was $1/\sqrt{512}$, on Text Classification the input sequence used is $1024$ and so the scale we used at initialization was $1/\sqrt{1024}$, on Retrieval the input sequence used is $2048$ and the initialized scale we used was $1/\sqrt{2048}$, on Image Classification the input sequence is $4096$ and the initialized scale we used was $1/\sqrt{4096}$, and on Pathfinder the input sequence used is $8192$ and the initialized scale we used was $1/\sqrt{8192}$. We followed the training regime of the original paper [30]. The results are shown in Table 4. As can be seen from the table the $x^{3}$ + dynamic scale activation results in better performance compared to softmax with $x^{3}$ + fixed scale performing comparably. However $x^{3}$ (with no scale) underperforms on all tasks.
Table 4: Comparisons of Nyströmformer models with different activation functions on the LRA benchmark. We report the accuracy (%).
ListOps Text Retrieval Image Pathfinder
@@ 6.3 Nyströmformer on LRA benchmark
$x^{3}$ + fixed scale 37.2 63.7 80.0 39.9 72.9
$x^{3}$ + dynamic scale 37.5 63.9 81.1 40.1 73.1
$x^{3}$ 32.3 62.0 78.5 38.1 67.9
7 Physics Informed Transformers
Recently, dynamical data has become essential for modeling time-dependent systems. [31] introduced transformer architectures for solving partial differential equations (PDEs), demonstrating superior performance across various PDE tasks. These models effectively capture interactions between spatial and temporal components inherent in dynamical systems. In Section A.2.5, we extend our polynomial attention framework to Physics-Informed Transformers, showing that polynomial activations can outperform softmax in this setting.
@@ 8 Limitations
While our work introduces polynomial activations that challenge the conventional understanding of softmax, there are several limitations to address. First, our theoretical framework is specifically developed for dot-product self-attention and may not directly generalize to other forms of attention mechanisms, such as additive attention or kernelized approximations. Extending our analysis to these variants could reveal additional nuances. Second, while our empirical evaluations span multiple architectures and tasks, they are limited to models with up to 100 million parameters due to resource constraints. The scalability of polynomial activations in large-scale transformers, especially those with billions of parameters, should be considered in the future.
@@ 9 Conclusion
This work questioned whether transformer activations for attention must produce sparse probability distributions. We introduced a theoretical framework analyzing the Frobenius norm of the self-attention matrix, which suggests key scaling properties for activations in attention mechanisms. We proved that specific polynomial activations, which behave very differently from softmax, satisfy these properties. Through extensive experiments across a variety of transformer applications, we demonstrated that these alternative activations not only compete with but can outperform softmax, offering a new perspective on attention mechanisms in transformers.
@@ A.1 Theoretical analysis
$\mathbf{softmax}\bigg{(}\begin{bmatrix}a_{11}&\cdots&a_{1n}\\ \vdots&\vdots&\vdots\\ a_{n1}&\cdots&a_{nn}\end{bmatrix}\bigg{)}=\begin{bmatrix}\frac{e^{a_{11}}}{\sum_{j=1}^{n}e^{a_{1j}}}&\cdots&\frac{e^{a_{1n}}}{\sum_{j=1}^{n}e^{a_{1j}}}\\ \vdots&\vdots&\vdots\\ \frac{e^{a_{n1}}}{\sum_{j=1}^{n}e^{a_{nj}}}&\cdots&\frac{e^{a_{nn}}}{\sum_{j=1}^{n}e^{a_{nj}}}.\end{bmatrix}$ (12)
@@
$||\mathbf{softmax}(A)||_{F}^{2}$ $=\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{1j}}}\bigg{)}^{2}(e^{2a_{11}}+\cdots e^{2a_{11}})+\cdots$ (13)
$\hskip 28.45274pt+\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{Nj}}}\bigg{)}^{2}(e^{2a_{N1}}+\cdots e^{2a_{NN}})$ (14)
$\leq\bigg{[}\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{1j}}}\bigg{)}(e^{a_{11}}+\cdots e^{a_{11}})\bigg{]}^{2}+\cdots$ (15)
$\hskip 28.45274pt+\bigg{[}\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{Nj}}}\bigg{)}(e^{a_{N1}}+\cdots e^{a_{NN}})\bigg{]}^{2}$ (16)
$=1+\cdots+1$ (17)
$=N$ (18)
@@
We move on to prove the second inequality in the statement of theorem 4.1. For this, let us write each entry of the matrix on the right of (12) as follows:
$F_{kl}=\frac{e^{a_{kl}}}{\sum_{j=1}^{N}e^{a_{kj}}}.$ (19)
@@
$\frac{\partial}{\partial x_{ij}}F_{ij}$ $=F_{ij}-F_{ij}^{2}$ (20)
$\frac{\partial}{\partial x_{ik}}F_{ij}$ $=-F_{ij}F_{ik}\text{ for any }k\neq j$ (21)
$\frac{\partial}{\partial x_{kl}}F_{ij}$ $=0\text{ for any }k\neq i\text{ and }l\neq j.$ (22)
@@
$\nabla\mathbf{softmax}(A)=\begin{bmatrix}\nabla F_{11}\\ \vdots\\ \nabla F_{1N}\\ \nabla F_{21}\\ \vdots\\ \vdots\\ \nabla F_{NN}\end{bmatrix}$ (23)
@@
$\nabla F_{ij}=\begin{bmatrix}-F_{ij}F_{i1}^{2}&-F_{ij}F_{i2}&\cdots&F_{ij}-F_{ij}^{2}&\cdots&F_{ij}F_{iN}.\end{bmatrix}$ (24)
@@
$||\nabla\mathbf{softmax}(A)||_{F}^{2}=||\nabla F_{11}||_{F}^{2}+\cdots+||\nabla F_{NN}||_{F}^{2}.$ (25)
@@
$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $=|F_{i1}-F_{i1}^{2}|^{2}+|F_{i1}F_{i2}|^{2}+\cdots+|F_{i1}F_{iN}|^{2}$ (26)
$\hskip 14.22636pt+|F_{i2}F_{i1}|^{2}+|F_{i2}-F_{i2}^{2}|^{2}+\cdots+|F_{i2}F_{iN}|^{2}$ (27)
@@
$\hskip 14.22636pt+|F_{iN}F_{i1}|^{2}+|F_{iN}F_{i2}|^{2}+\cdots+|F_{iN}-F_{iN}^{2}|^{2}$ (29)
$\leq(F_{i1})^{2}(|1-F_{i1}|+|F_{i2}|+\cdots+|F_{iN}|)^{2}$ (30)
$\hskip 14.22636pt(F_{i2})^{2}(|F_{i1}|+|1-F_{i2}|+\cdots+|F_{iN}|)^{2}$ (31)
@@
$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $\leq 4F_{i1}^{2}(\widehat{F_{i1}}+F_{i2}+\cdots+F_{iN})$ (34)
@@
$\hskip 14.22636pt+4F_{iN}^{2}(F_{i1}+F_{i2}+\cdots+\widehat{F_{iN}})$ (36)
$\leq 4(F_{i1}^{2}+\cdots F_{iN}^{2})$ (37)
$=4.$ (38)
@@
$||\nabla\mathbf{softmax}(A)||_{F}^{2}\leq 4N$ (39)
@@
$||\nabla\mathbf{softmax}(A)||_{F}\leq 2\sqrt{N}.$ (40)
@@
$\mathop{\mathbb{E}}(c_{11}^{2})$ $=\mathop{\mathbb{E}}((\sum_{i=1}^{D}a_{1i}b_{i1})^{2})$ (41)
@@
$\mathop{\mathbb{E}}(c_{11}^{2})$ $=\mathop{\mathbb{E}}((\sum_{i=1}^{D}a_{1i}b_{i1})^{2p})$ (42)